the sum of the ages of Lois and Leonard is 29years. Lois is 3 years than Leonard.
how old are they?

Answer by ikleyn(52779)

.
the sum of the ages of Lois and Leonard is 29years. Lois is 3 years than Leonard.
how old are they?
~~~~~~~~~~~~~~~~~~~~~


"Lois is 3 years . . . W H A T   ?    ? ?     ? ? ? . . . than Leonard"


        - and you failed to provide an exact, correct and precise input.


------------
Comment from student: The sum of the ages of Lois and Leonard is 29 years. Lois is 3 years older than Leonard.
how old are they please this is the correct question sorry for the mistake
------------


My response :

x + y = 29 (1)
x - y = 3 (2)
------------------------- Add the equations

2x = 29 + 3
2x = 32 =============> x = 32/2 = 16 years is the age of the older person.

Then y = 29 - 16 = 13 years is the age of the younger person.

Solved.